Find two consecutive even integers such that two times the first plus three times the 2nd is 76.
Answer by checkley71(8403)

2(x)+3(x+2)=76
2x+3x+6=76
5x=76-6
5x=70
x=70/5
x=14 first even number.
14+2=16 next even number.
proof
2*14+3*16=76
28+48=76
76=76